# Fast and scalable Wasserstein-1 neural optimal transport solver for single-cell perturbation prediction

### v1

@@ Abstract
Predicting single-cell perturbation responses requires mapping between two unpaired single-cell data distributions. Optimal transport (OT) theory provides a principled framework for constructing such mappings by minimizing transport cost. Recently, Wasserstein-2 ($W_{2}$) neural optimal transport solvers (e.g., CellOT) have been employed for this prediction task. However, $W_{2}$ OT relies on the general Kantorovich dual formulation, which involves optimizing over two conjugate functions, leading to a complex min-max optimization problem that converges slowly. To address these challenges, we propose a novel solver based on the Wasserstein-1 ($W_{1}$) dual formulation. Unlike $W_{2}$, the $W_{1}$ dual simplifies the optimization to a maximization problem over a single 1-Lipschitz function, thus eliminating the need for time-consuming min-max optimization. While solving the $W_{1}$ dual only reveals the transport direction and does not directly provide a unique optimal transport map, we incorporate an additional step using adversarial training to determine an appropriate transport step size, effectively recovering the transport map. Our experiments demonstrate that the proposed $W_{1}$ neural optimal transport solver can mimic the $W_{2}$ OT solvers in finding a unique and “monotonic” map on 2D datasets. Moreover, the $W_{1}$ OT solver achieves performance on par with or surpasses $W_{2}$ OT solvers on real single-cell perturbation datasets. Furthermore, we show that $W_{1}$ OT solver achieves 2545×25$\sim$ 45$\times$25 ∼ 45 × speedup, scales better on high dimensional transportation task, and can be directly applied on single-cell RNA-seq dataset with highly variable genes. Our implementation and experiments are open-sourced at https://github.com/poseidonchan/w1ot.
@@
Optimal transport Wasserstein-1 Lipschitz neural network Single-cell data
@@ 1 Introduction
Single-cell perturbation experiments have emerged as a powerful tool for understanding cellular responses to environmental changes, such as drug treatments and other stimuli. These experiments are crucial for characterizing how cells react to perturbations, which has significant implications for fields such as cancer research [1] and drug development [2]. Common approaches for single-cell perturbation experiments include single-cell RNA sequencing (scRNA-seq) [3] and Iterative Indirect Immunofluorescence Imaging (4i) [4]. scRNA-seq allows for the measurement of gene expression changes in response to genetic [5] or drug [6] perturbations, while 4i provides spatially resolved information on protein abundance and localization.
@@
Several methods have been proposed to address the challenge of predicting single-cell perturbation responses. The baseline method, scGen [7], calculates the shift between the means of embedded target and source cells and applies it uniformly to transport the data. Building on this idea of applying a shift in the latent space, other methods aim to learn disentangled representations to model perturbation responses. For instance, chemCPA [8, 9] employs an encoder-decoder structure to learn disentangled representations of cells and drugs, allowing for additive operations in the latent space to predict perturbation effects. Similarly, biolord [10] uses latent optimization techniques to better model different attributes. Although these methods can capture nonlinear responses, they transform the perturbation prediction problem into the potentially more challenging task of learning invariant and disentangled representations.
Recently, computational methods based on optimal transport (OT) theory — such as CellOT [11] and CINEMA-OT [12] — have offered a principled framework for mapping between unpaired distributions. However, CINEMA-OT is designed to model existing perturbation data with a discrete OT formulation and thus cannot be used for prediction tasks. On the other hand, CellOT is built on the continuous Wasserstein-2 ($W_{2}$) OT formulation [13] and can be used for generative prediction. Nevertheless, the $W_{2}$ formulation requires optimizing a complex min-max problem and optimizing input convex neural networks (ICNNs) [14], which leads to slow convergence and poor performance on high-dimensional data, respectively [15].
To address the computational and scalability limitations of existing $W_{2}$ OT solvers, we propose a novel solver based on the Wasserstein-1 ($W_{1}$) formulation. This approach offers two key advantages. First, the $W_{1}$ dual simplifies the min-max optimization problem over two convex functions into a maximization problem over a single 1-Lipschitz function [16, 17], thereby avoiding the complexities inherent in min-max optimization. Second, by eliminating the need to parameterize convex functions using ICNNs, our approach achieves better scalability, as ICNNs can be difficult to optimize in high-dimensional settings [15].
In this paper, we develop the first accurate $W_{1}$ OT solver leveraging recent advancements in 1-Lipschitz neural networks [18, 19] and generative adversarial networks (GANs) [20]. We demonstrate that our solver is capable of 1) finding the “monotonic” transport map on 2 dimensional (2D) datasets; 2) achieving similar or better performance on real single-cell perturbation datasets; 3) being 2545×25$\sim$ 45$\times$25 ∼ 45 × faster and more scalable on high dimensional datasets.
2 Results
@@ 2.1 Method overview
Figure 1: Overview of $W_{1}$
@@
The single-cell perturbation task requires finding an unknown underlying map to transform the distribution of control group cells ($\mu$) to the distribution of perturbation group cells ($\nu$) in a high-dimensional space (Fig. 1a). Most current single-cell measurements destroy the cells, resulting in unpaired single-cell perturbation datasets. Optimal transport is a principled approach to map these unpaired data distributions. It assumes cells change following a “minimal effort” principle, which has proven effective in many single-cell modeling problems [21]. Previous generative OT map models the transport effort as quadratic cost and has limitations on optimization complexity and scalability [13, 11, 15]. Since these limitations arise from the use of the $W_{2}$ Kantorovich dual, we instead develop an OT solver based on the $W_{1}$ dual (also known as the Kantorovich-Rubinstein dual) to avoid the min-max problem [16]. However, building a $W_{1}$ OT solver presents theoretical challenges, as solving the dual problem does not directly recover the primal problem of finding the OT map; instead, it only provides insight into the transportation direction [22, 23]. Based on these analyses, our proposed $W_{1}$ OT solver learning procedure is divided into two steps: it first learns the transport direction $\nabla f$ by maximizing the $W_{1}$ dual, and then it learns the transport step size $\eta$ based on the transport direction guidance (Fig. 1b). We parameterize the 1-Lipschitz Kantorovich potential $f$ as a GroupSort neural network [18] and recover the transport direction with its gradient $\nabla f$. After fixing the transport direction, we parameterize the step size function $\eta$ as a non-negative deep neural network (DNN) and employ an additional DNN discriminator to adversarially train it to learn the appropriate step size. More background and details are included in the Methods section.
2.2 $W_{1}$ OT solver learns the monotonic map on 2D datasets
Figure 2: $W_{1}$ solver on toy datasets
Besides the challenge of determining the transport step size in $W_{1}$ OT, another challenge arises from the fact that we may construct multiple transport plans solely based on the transport direction guidance. Non-uniqueness occurs in one-dimensional transport scenarios when the data points in the source and target distributions are collinear. This may cause the optimal transport solution to lose the desired “monotonicity” property [23, 24]. In this context, “monotonicity” refers to preserving the relative order of points after transport, as discussed in the literature. In the case of mapping two cell distributions, “monotonicity” can be understood as preserving the local cell type structure after transport (see details in Methods).
According to previous theoretical analyses, as long as the constructed transport map satisfies a certain condition (see Equation 6 in Methods), the transport map is unique and exhibits “monotonicity”. However, this condition is intractable because it requires evaluating every pair of data points in the source distribution. Therefore, it remains a question of whether our proposed $W_{1}$ OT solver can find the desired “monotonic” transport map and remain biologically meaningful in this application scenario.
To validate that our solver can learn the “monotonic” optimal transport map, we begin by evaluating our method on two specially designed 2D datasets (see Methods). First, we test the collinear 1D scenario, referred to as the “bookshelf,” where the source and target distributions are segments of the same length on the x-axis (Fig. 2a). After learning the transport direction, it is possible to construct more than one transport map with the same optimal cost. For instance, one solution could involve moving the rightmost point in the source distribution to the leftmost point in the target distribution, thus inverting the order of the entire segment. Another solution would involve shifting all points in the source distribution to the target distribution while maintaining the original order. In the experiment, we place special markers in the source distribution and observe that the relative order of the transported markers is preserved (Fig. 2a). Moreover, we design a 2D example consisting of concentric circles to demonstrate that our solver also learns the “monotonic” map in a 2D scenario (Fig. 2b). In this example, the inner source distribution is expected to map to the inner target distribution, rather than the outer target distribution. If not, the transport map would lose its “monotonicity” as the inner-outer relative order would be disrupted. As we can see, the learned transport map faithfully recovers the “monotonic” map, since there are no transport rays connecting the outer source distribution to the inner target distribution (Fig. 2b). Next, we evaluate the $W_{1}$ OT solver on two difficult non-linear examples named “swiss roll” and “moons” (Fig. 2c,d). The results show that our solver can successfully learn the non-trivial transport map and exhibit similar performance to the $W_{2}$ OT solver in literature [25].
2.3 $W_{1}$ OT solver accurately predicts single-cell responses
Figure 3: Performance benchmark on 2 real single-cell datasets.
We apply the $W_{1}$ OT solver to real single-cell perturbation datasets to verify its effectiveness. We also benchmark other single-cell perturbation prediction methods, including $W_{2}$ OT solver [13, 11] and scGen [7]. Since our primary objective is to develop a fast and scalable OT solver, we exclude chemCPA [9] and biolord [10] from the experiments. We also report the results of directly using control group cells or perturbation group cells as prediction outputs, labeled as identity and observed, respectively, to demonstrate the lower and upper bound performance. We conduct our experiments on two datasets: the “4i” imaging dataset [11] and the “sciplex3” scRNA-seq dataset [6]. The “4i” imaging dataset contains 48 features for 37 perturbations while the “sciplex3” scRNA-seq dataset contains 188 perturbations and 1,000 highly variable genes after selection. To reduce the experiments overhead, we follow the experiments in chemCPA [9] to test 9 drugs: Dacinostat, Givinostat, Belinostat, Hesperadin, Quisinostat, Alvespimycin, Tanespimycin, TAK-901, and Flavopiridol, as these drugs were reported among the most effective drugs in the original publication [6]. For the low dimensional imaging dataset, both OT solvers are directly applied to this original feature space. For the high-dimensional scRNA-seq dataset, both OT solvers are applied to the latent space (50 features) of an autoencoder and then decoded back to cells. Each method is independently run 5 times with randomly splitted training and testing sets.
To evaluate performance, we use three metrics: $r^{2}$ (correlation coefficient), $l_{2}$ (Euclidean distance) between feature means, and Maximum Mean Discrepancy (MMD) [26]. While $r^{2}$ and $l_{2}$ only assess differences between the means of the transported and target cell populations, MMD provides a more comprehensive evaluation by measuring the alignment of the entire distribution. For the high dimensional scRNA-seq dataset, we evaluate the top 50 differentially expressed genes between the control and perturbation groups to avoid the bias [11].
We report the performance of each method on both datasets by summarizing the results across all perturbations (Fig. 3a,b). We use the Wilcoxon rank-sum test to compare the performance differences between the $W_{1}$ OT solver and the two baselines. The results show that the $W_{1}$ OT solver generally achieves statistically similar performance to the $W_{2}$ OT solver and significantly outperforms scGen on both datasets. Furthermore, we observe that the performance of both OT solvers closely matches the observed target distribution, demonstrating their effectiveness and accuracy. Interestingly, we find that while scGen performs well on the $r^{2}$ and $l_{2}$ metrics, it is less competitive on the MMD metric. This can be attributed to the fact that scGen primarily focuses on aligning population means but neglects the alignment of the overall distribution. This observation underscores the importance of using optimal transport to map the entire distribution to the target, rather than just aligning the means.
Figure 4: $W_{1}$ solver on high dimensional datasets.
We further test the scalability of the $W_{1}$ OT solver by measuring its training time and performance on high-dimensional datasets. To demonstrate the efficiency and speed of the $W_{1}$ OT solver, we compare its training time to that of the $W_{2}$ OT solver across various data dimensions, using an 8-core Intel Core i9-11900K processor (Fig. 4a). The results show that the $W_{1}$ OT solver is approximately 2.5 to 4.5 times faster than the $W_{2}$ OT solver for the same number of training iterations. Given that the default training iteration for $W_{2}$ OT is set at 100,000 in different studies [13, 11], the overall speedup of $W_{1}$ OT would be in the range of 25 to 45 times. In practical terms, while the $W_{2}$ OT solver requires about an hour and a half to train, our proposed $W_{1}$ OT solver can complete training in just 5 minutes on a CPU. This speedup would significantly accelerate single-cell perturbation research and make training on ultra-large-scale datasets feasible.
To validate that the performance of the $W_{1}$ OT solver is also scalable, we applied both OT solvers directly to the “sciplex3” dataset with 1000 features. We observe that the performance of the $W_{1}$ OT solver on high-dimensional data remains consistent with its performance on the latent space of 50 features (Fig. 3 and Fig. 4b). However, the performance of the $W_{2}$ OT solver drops significantly, even falling below the lower bound (identity transformation). This indicates that solving the min-max problem is highly challenging and impractical in high-dimensional settings which has also been confirmed by previous studies [15].
3 Discussions
In this work, we present a novel optimal transport solver based on the Wasserstein-1 dual formulation, which offers significant computational advantages over Wasserstein-2 neural optimal transport solvers [13, 15, 25]. Our approach leverages 1-Lipschitz GroupSort networks combined with GANs to construct an efficient and scalable solver. Through extensive experiments, we demonstrate that our solver learns biologically meaningful “monotonic” transport maps and achieves a 2545×25$\sim$ 45$\times$25 ∼ 45 × speedup while maintaining comparable performance to existing $W_{2}$ OT methods on real single-cell datasets.
A key advantage of our $W_{1}$ OT solver is its scalability for high-dimensional data, which stems from the use of the $W_{1}$ dual formulation and 1-Lipschitz GroupSort networks. Although previous work questions the capability of 1-Lipschitz GroupSort networks to accurately estimate $W_{1}$ distance in high-dimensional settings [27], we demonstrate that these networks perform reasonably well on high-dimensional scRNA-seq datasets when implemented with appropriate grouping sizes [19]. This capability is crucial as the field advances toward high-dimensional universal cell representations using single-cell foundation models [28, 29, 30].
Our approach, which constructs 1-Lipschitz neural networks layer by layer, can be readily extended to model conditional transport maps $T(x,\text{condition})$. Similar to how partially input convex neural networks (PICNN) [14] can be used to build conditional OT maps in the $W_{2}$ setup [31], we can construct a partially 1-Lipschitz neural network layer $h$ as follows: $h(x,\text{condition})=f(x)+g(\text{condition})$, where $f$ is 1-Lipschitz. With this construction, we can incorporate different conditions in the optimal transport map and enable efficient transport mapping for complex real-world applications [31].
While we empirically validate that the $W_{1}$ OT solver can learn “monotonic” maps on 2D datasets, there is currently no theoretical explanation or guarantee for this phenomenon. Additionally, although we demonstrate our solver’s strong performance in various scenarios, GroupSort neural networks have not been proven to be universal 1-Lipschitz approximators under the Euclidean norm [18], which may limit their performance in certain cases.
4 Methods
@@ 4.1.1 Wasserstein-1 optimal transport
In optimal transport, we aim to find the optimal transport plan to transport one distribution to another in a sense of minimum total cost. If the cost is defined as the conventional distance metric which satisfies the triangular inequality, it is called Wasserstein-1 optimal transport. If the cost is defined as the distance metric to the power of $p$, then the problem is called as the Wasserstein-$p$ optimal transport. Compared to the Wasserstein-$p$ metric, the Wasserstein-1 optimal transport will be more robust to the outlier and the noise, and its elegant dual form will be numerically easy to solve [17]. Hence, we focus on discussing the Wasserstein-1 optimal transport and its practical solver below.
@@
where $\mathrm{\Phi}_{c}={(\phi,\psi):\phi(y)-\psi(x)\leq c(x,y),\forall x\in\mathcal{X},y\in\mathcal{Y}}$. To achieve the supremum (i.e., optimal cost or the Wasserstein distance), the functions $\psi$ and $\phi$ are related by the $c$-transform: $\phi(y)=\psi^{c}(y)=\inf_{x}(c(x,y)+\psi(x))$. When the cost function is the distance metric, the Wasserstein-1 distance $W_{1}(\mu,\nu)$ can be expressed using the Kantorovich-Rubinstein duality formula [16]:
@@
where $f$ is constrained to be a 1-Lipschitz continuous function and is referred as Kantorovich potential function in literature. Notably, the dual formulation of the Wasserstein-1 distance significantly simplifies the optimization procedure since we only need to optimize over a single function. In contrast, solving for the Wasserstein-$p$ ($p>1$) is generally more challenging and often involves a min-max optimization problem [13, 11].
@@
Unlike the Wasserstein-$p$ ($p>1$) scenario, where the optimal transport map can be uniquely determined by the gradient of the Kantorovich potential (i.e., Brenier’s theorem [32]) due to the strictly convex cost function, Wasserstein-1 Kantorovich potential only reveals the transport direction and its optimal transport has non-unique solutions due to the undetermined step size in the general setting [33, 23, 22]:
@@
For example, imagine we have $n$ identical books lined up on a bookshelf, occupying positions 1 through $n$. If we want to optimally transport these books so they occupy positions 2 through $n+1$, there are multiple valid ways to do this. One solution could be to move the first book directly to position $n+1$, while another could be to shift each book one position to the right. Both solutions have the same total cost of $n$. We can see that in the Wasserstein-1 optimal transport setting, we only know the optimal transport direction but lack information on the transported distance for each point, which gives us multiple options to construct the final optimal transport map. Moreover, previous studies have shown that the non-uniqueness 1-D phenomenon is the only source of non-uniqueness in Monge’s problem [24].
@@
Among the various possible optimal transport maps, researchers aim to construct one that mimics the Wasserstein-$p$ optimal transport map and preserves the desired “monotonicity” property. Returning to the previously introduced 1-D example, we might prefer the second optimal transport map because it “monotonically” shifts all books to the next position, thereby preserving the local structure. Similarly, when mapping one single-cell data distribution to another, the desired transport is expected to maintain the local structure (e.g., cell type geometric structure) to preserve the biological meaning. Theoretically, previous geometric analysis has shown that this map exists, is unique, and should satisfy the following criterion [23, 24]:
@@
where $x_{1}\neq x_{2}\in\mathcal{X}$ and $T(x_{1})\neq T(x_{2})$. Furthermore, it has been shown that the inversion of (5) occurs only when $x_{1},x_{2},T(x_{1}),T(x_{2})$ are collinear (i.e., $x_{1}$ and $x_{2}$ share the same transport direction, ${\nabla f(x_{1})}={\nabla f(x_{2})}$) [24]. This condition indicates that when they are collinear, the direction of the vector $x_{1}-x_{2}$ is the same as the direction of the vector $T(x_{1})-T(x_{2})$. In other words, the transported points should maintain their relative order as in the source distribution.

### v2

@@ Abstract
Motivation: Predicting single-cell perturbation responses requires mapping between two unpaired single-cell data distributions. Optimal transport (OT) theory provides a principled framework for constructing such mappings by minimizing transport cost. Recently, Wasserstein-2 ($W_{2}$) neural optimal transport solvers (e.g., CellOT) have been employed for this prediction task. However, $W_{2}$ OT relies on the general Kantorovich dual formulation, which involves optimizing over two conjugate functions, leading to a complex min-max optimization problem that converges slowly.
Results: To address these challenges, we propose a novel solver based on the Wasserstein-1 ($W_{1}$) dual formulation. Unlike $W_{2}$, the $W_{1}$ dual simplifies the optimization to a maximization problem over a single 1-Lipschitz function, thus eliminating the need for time-consuming min-max optimization. While solving the $W_{1}$ dual only reveals the transport direction and does not directly provide a unique optimal transport map, we incorporate an additional step using adversarial training to determine an appropriate transport step size, effectively recovering the transport map. Our experiments demonstrate that the proposed $W_{1}$ neural optimal transport solver can mimic the $W_{2}$ OT solvers in finding a unique and “monotonic” map on 2D datasets. Moreover, the $W_{1}$ OT solver achieves performance on par with or surpasses $W_{2}$ OT solvers on real single-cell perturbation datasets. Furthermore, we show that $W_{1}$ OT solver achieves 2545×25$\sim$ 45$\times$25 ∼ 45 × speedup, scales better on high dimensional transportation task, and can be directly applied on single-cell RNA-seq dataset with highly variable genes.
Availability and Implementation: Our implementation and experiments are open-sourced at https://github.com/poseidonchan/w1ot.
Contact: Yanshuo Chen (cys@umd.edu) and Heng Huang (heng@umd.edu).
Supplementary Information: Supplementary data are available online at the journal’s website.
@@
Perturbation, Single-cell data, Wasserstein-1 optimal transport
@@ 1 Introduction
Single-cell perturbation experiments have emerged as a powerful tool for modeling cellular responses to environmental changes, such as drug treatments and other stimuli. These experiments are crucial for characterizing how cells react to perturbations, which has significant implications for fields such as cancer research (Frangieh et al., 2021) and drug development (Van De Sande et al., 2023). Common approaches for single-cell perturbation experiments include single-cell RNA sequencing (scRNA-seq) (Peidli et al., 2024) and Iterative Indirect Immunofluorescence Imaging (4i) (Gut et al., 2018). scRNA-seq allows for the measurement of gene expression changes in response to genetic (Norman et al., 2019) or drug (Srivatsan et al., 2020) perturbations, while 4i provides spatially resolved information on protein abundance and localization.
@@
Several methods have been proposed to address the challenge of predicting single-cell perturbation responses. The baseline method, scGen (Lotfollahi et al., 2019), calculates the shift between the means of embedded target and source cells and applies it uniformly to transport the data. Building on this idea of applying a shift in the latent space, other methods aim to learn disentangled representations to model perturbation responses. For instance, chemCPA (Lotfollahi et al., 2023; Hetzel et al., 2022) employs an encoder-decoder structure to learn disentangled representations of cells and drugs, allowing for additive operations in the latent space to predict perturbation effects. Similarly, biolord (Piran et al., 2024) uses latent optimization techniques to better model different attributes. Although these methods can capture nonlinear responses, they transform the perturbation prediction problem into the potentially more challenging task of learning invariant and disentangled representations.
Recently, computational methods based on optimal transport (OT) theory — such as CellOT (Bunne et al., 2023) and CINEMA-OT (Dong et al., 2023) — have offered a principled framework for mapping between unpaired distributions. However, CINEMA-OT is designed to model existing perturbation data with a discrete OT formulation and thus cannot be used for prediction tasks. On the other hand, CellOT is built on the continuous Wasserstein-2 ($W_{2}$) OT formulation (Makkuva et al., 2020) and can be used for generative prediction. Nevertheless, the $W_{2}$ formulation requires optimizing a complex min-max problem and optimizing input convex neural networks (ICNNs) (Amos et al., 2017), which leads to slow convergence and poor performance on high-dimensional data, respectively (Korotin et al., 2021).
To address the computational and scalability limitations of existing $W_{2}$ OT solvers, we propose a novel solver based on the Wasserstein-1 ($W_{1}$) formulation. This approach offers two key advantages. First, the $W_{1}$ dual simplifies the min-max optimization problem over two convex functions into a maximization problem over a single 1-Lipschitz function (Villani et al., 2009; Peyré et al., 2019), thereby avoiding the complexities inherent in min-max optimization. Second, by eliminating the need to parameterize convex functions using ICNNs, our approach achieves better scalability, as ICNNs can be difficult to optimize in high-dimensional settings (Korotin et al., 2021).
In this paper, we develop the first accurate $W_{1}$ OT solver leveraging recent advancements in 1-Lipschitz neural networks (Anil et al., 2019; Tanielian and Biau, 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014). We demonstrate that our solver is capable of 1) finding the “monotonic” transport map on 2 dimensional (2D) datasets; 2) achieving similar or better performance on real single-cell perturbation datasets; 3) being 2545×25$\sim$ 45$\times$25 ∼ 45 × faster and more scalable on high dimensional datasets.
2 Materials and methods
@@ 2.1 Method overview
Figure 1: Overview of $W_{1}$ a. The single-cell perturbation prediction task needs a mapping to map the control group ($\mu$) cells to the perturbation group ($\nu$) cells. b. Our proposed $W_{1}$ OT solver. The first step is to learn the transport direction by maximizing the Kantorovich dual problem. The second step is to learn the appropriate transport step size via adversarial training to finally construct the transport map.Figure conceptually inspired by the illustration in the CellOT paper Bunne et al. (2023).
@@
The single-cell perturbation task requires finding an unknown underlying map to transform the distribution of control group cells ($\mu$) to the distribution of perturbation group cells ($\nu$) in a high-dimensional space (Fig. 1a). Most current single-cell measurements destroy the cells, resulting in unpaired single-cell perturbation datasets. Optimal transport is a principled approach to map these unpaired data distributions. It assumes cells change following a “minimal effort” principle, which has proven effective in many single-cell modeling problems (Bunne et al., 2024). Previous generative OT map models the transport effort as quadratic cost and has limitations on optimization complexity and scalability (Makkuva et al., 2020; Bunne et al., 2023; Korotin et al., 2021). Since these limitations arise from the use of the $W_{2}$ Kantorovich dual, we instead develop an OT solver based on the $W_{1}$ dual (also known as the Kantorovich-Rubinstein dual) to avoid the min-max problem (Villani et al., 2009). However, building a $W_{1}$ OT solver presents theoretical challenges, as solving the dual problem does not directly recover the primal problem of finding the OT map; instead, it only provides insight into the transportation direction (Trudinger and Wang, 2001; Caffarelli et al., 2002). Based on these analyses, our proposed $W_{1}$ OT solver learning procedure is divided into two steps: it first learns the transport direction $\nabla f$ by maximizing the $W_{1}$ dual, and then it learns the transport step size $\eta$ based on the transport direction guidance (Fig. 1b). We parameterize the 1-Lipschitz Kantorovich potential $f$ as a GroupSort neural network (Anil et al., 2019) and recover the transport direction with its gradient $\nabla f$. After fixing the transport direction, we parameterize the step size function $\eta$ as a non-negative deep neural network (DNN) and employ an additional DNN discriminator to adversarially train it to learn the appropriate step size. More background and details are included below.
@@ 2.2.1 Wasserstein-1 optimal transport
In optimal transport, we aim to find the optimal transport plan to transport one distribution to another in a sense of minimum total cost. If the cost is defined as the conventional distance metric which satisfies the triangular inequality, it is called Wasserstein-1 optimal transport. If the cost is defined as the distance metric to the power of $p$, then the problem is called as the Wasserstein-$p$ optimal transport. Compared to the Wasserstein-$p$ metric, the Wasserstein-1 optimal transport will be more robust to the outlier and the noise, and its elegant dual form will be numerically easy to solve (Peyré et al., 2019). Hence, we focus on discussing the Wasserstein-1 optimal transport and its practical solver below.
@@
where $\mathrm{\Phi}_{c}={(\phi,\psi):\phi(y)-\psi(x)\leq c(x,y),\forall x\in\mathcal{X},y\in\mathcal{Y}}$. To achieve the supremum (i.e., optimal cost or the Wasserstein distance), the functions $\psi$ and $\phi$ are related by the $c$-transform: $\phi(y)=\psi^{c}(y)=\inf_{x}(c(x,y)+\psi(x))$. When the cost function is the distance metric, the Wasserstein-1 distance $W_{1}(\mu,\nu)$ can be expressed using the Kantorovich-Rubinstein duality formula (Villani et al., 2009):
@@
where $f$ is constrained to be a 1-Lipschitz continuous function and is referred as Kantorovich potential function in literature. Notably, the dual formulation of the Wasserstein-1 distance significantly simplifies the optimization procedure since we only need to optimize over a single function. In contrast, solving for the Wasserstein-$p$ ($p>1$) is generally more challenging and often involves a min-max optimization problem (Makkuva et al., 2020; Bunne et al., 2023).
@@
Unlike the Wasserstein-$p$ ($p>1$) scenario, where the optimal transport map can be uniquely determined by the gradient of the Kantorovich potential (i.e., Brenier’s theorem (Brenier, 1991)) due to the strictly convex cost function, Wasserstein-1 Kantorovich potential only reveals the transport direction and its optimal transport has non-unique solutions due to the undetermined step size in the general setting (Gangbo and McCann, 1996; Caffarelli et al., 2002; Trudinger and Wang, 2001):
@@
For example, imagine we have $n$ identical books lined up on a bookshelf, occupying positions 1 through $n$. If we want to optimally transport these books so they occupy positions 2 through $n+1$, there are multiple valid ways to do this. One solution could be to move the first book directly to position $n+1$, while another could be to shift each book one position to the right. Both solutions have the same total cost of $n$. We can see that in the Wasserstein-1 optimal transport setting, we only know the optimal transport direction but lack information on the transported distance for each point, which gives us multiple options to construct the final optimal transport map. Moreover, previous studies have shown that the non-uniqueness 1-D phenomenon is the only source of non-uniqueness in Monge’s problem (Feldman and McCann, 2002).
@@
Among the various possible optimal transport maps, researchers aim to construct one that mimics the Wasserstein-$p$ optimal transport map and preserves the desired “monotonicity” property. Returning to the previously introduced 1-D example, we might prefer the second optimal transport map because it “monotonically” shifts all books to the next position, thereby preserving the local structure. Similarly, when mapping one single-cell data distribution to another, the desired transport is expected to maintain the local structure (e.g., cell type geometric structure) to preserve the biological meaning. Theoretically, previous geometric analysis has shown that this map exists, is unique, and should satisfy the following criterion (Caffarelli et al., 2002; Feldman and McCann, 2002):
@@
where $x_{1}\neq x_{2}\in\mathcal{X}$ and $T(x_{1})\neq T(x_{2})$. Furthermore, it has been shown that the inversion of (6) occurs only when $x_{1},x_{2},T(x_{1}),T(x_{2})$ are collinear (i.e., $x_{1}$ and $x_{2}$ share the same transport direction, ${\nabla f(x_{1})}={\nabla f(x_{2})}$) (Feldman and McCann, 2002). This condition indicates that when they are collinear, the direction of the vector $x_{1}-x_{2}$ is the same as the direction of the vector $T(x_{1})-T(x_{2})$. In other words, the transported points should maintain their relative order as in the source distribution.